OpenStudy (jordanloveangel):

2)What is the solution to the system of equations? x+3y +5 x-4y=3 (___ , ___)

OpenStudy (jordanloveangel):

@Clonetroooper78

OpenStudy (jordanloveangel):

@TheEdwardsFamily

OpenStudy (theedwardsfamily):

dont know this

OpenStudy (jordanloveangel):

k

OpenStudy (jordanloveangel):

@fallenangelorchid

OpenStudy (jordanloveangel):

@cody77445

OpenStudy (anonymous):

is there an = sign in the first equation

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (jordanloveangel):

hold on let me see

OpenStudy (jordanloveangel):

@cody77445

OpenStudy (anonymous):

ok that makes more sense

OpenStudy (jordanloveangel):

k

OpenStudy (anonymous):

i got x=11 y=2

OpenStudy (anonymous):

but i could be wrong

OpenStudy (jordanloveangel):

oh k thx can u help with few more please

OpenStudy (anonymous):

sure

OpenStudy (jordanloveangel):

Solve the given system, using the substitution method. y = 3x – 7 6x – 2y = 12 A. (14, 12) B. There are an infinite number of solutions. C. (12, 14) D. There is no solution.

OpenStudy (jordanloveangel):

i think its D

OpenStudy (anonymous):

me too btw are you in connexus

OpenStudy (jordanloveangel):

no im not

OpenStudy (anonymous):

dang

OpenStudy (jordanloveangel):

lol

OpenStudy (jordanloveangel):

Solve, using the substitution method. y – 3x = 7 21 + 9x = 3y A. There is no solution. B. There are an infinite number of solutions. C. The solution is (1, 10) D. The solution is (28, 0)

OpenStudy (jordanloveangel):

A?